# RiTeK: A Dataset for Large Language Models Complex Reasoning over Textual Knowledge Graphs

### v1

@@ 4.2.3 Data Distribution Analysis
We chose Shannon Entropy and Type-Token Ratio (TTR) as the metrics to evaluate the query diversity generated in our two datasets. Shannon Entropy takes into account the frequency of each word, measuring the evenness of word distribution in the text, while Type-Token Ratio reflects the variety of words, with a higher value indicating greater diversity in the generated queries. As shown in Table 4, the TTR values for both RiTeK-PharmKG and RiTeK-ADint surpass those of STARK-Prime, demonstrating that the queries generated in our datasets exhibit high complexity and diversity. For Shannon Entropy, our results are comparable to STARK-Prime. Since our RiTeK-ADint dataset involves a wide range of non-pharmacological interventions (NPIs), lifestyle modifications, and environmental factors, it introduces a richer variety of specialized terminology and concepts into the synthesized queries. This expanded vocabulary diversity leads to significantly higher Shannon Entropy compared to the other medical domain datasets. However, since our two datasets are derived from the medical domain, the frequent repetition of specialized medical terminology, as well as the more concentrated vocabulary compared to general-domain texts, results in slightly lower Shannon Entropy for our datasets than for the other two general-domain datasets. For more analysis about the distribution of query lengths and answer length, please refer Appendix A.3.
5 Proposed Approach: Relational MCTS
5.1 Motivation and Approach Overview
Monte Carlo Tree Search (MCTS) has made significant advancements in mathematics, as it dynamically explores and evaluates potential solutions, balancing exploration and exploitation to optimize decision-making in complex, high-dimensional spaces. However, the effectiveness of MCTS in TKGs has not yet been explored. To address this gap, this paper investigates the effectiveness of MCTS and proposes an improved version Relational MCTS that dynamically retrieves relational paths from the TKGs for the special query. Specifically, it involves four steps: 1) To improve the retrieval efficiency, we begin our path retrieval by creating the textual triple graph (TTG) for each textual KG. 2) Subsequently, the node embeddings in the TTG and the input query embedding are computed using a pre-trained language model, such as a SentenceBERT Baek et al. (2023). Notably, each node embedding includes its associated textual information. These embeddings will be used to calculate the reward in the MCTS and Relational MCTS. 3) Proposed Relation MCTS is used to search the relational path information. 4) Finally, the retrieved relational path information will be verbalized and fed into the LLM, such as GPT-4, along with the input query to predict the final answers. Please refer to Appendix B for detailed information on steps 2 and 3.
Shannon Entropy Type-Token Ratio
Medical domain
RiTeK-ADint 10.04 0.187
RiTeK-PharmKG 9.61 0.157
STARK-PRIME 9.63 0.143
General domain
STARK-AMAZON 10.39 0.179
STARK-MAG 10.25 0.180
Table 4: Query diversity measurement
5.2 MCTS vs Relational MCTS
In traditional MCTS, a trajectory typically extends from the root to a terminal node. However, in the context of QA tasks over textual knowledge graphs, many relational templates incorporate constraints, which pose challenges for standard MCTS. Furthermore, the depth of the reasoning path can vary depending on the specific question being posed; for example, some queries may require a reasoning depth of three hops, while others may only require two. Consequently, MCTS is hard to appropriately recognize or adapt to the required depth for each individual query. To address this issue, we propose an enhanced MCTS specifically designed to retrieve the relevant inference path information for the input query, which we refer to as Relational MCTS.
Figure 3: Expansion progress of MCTS and Relational MCTS. The pink shallow circle refers to the $Stop$ sign. Each node $tn$ refers to a triple.
Same as traditional MCTS, starting from the initial topic entity $e$, we perform multiple searches consisting of selection, expansion, simulations, and back-propagation. The key distinction between MCTS and Relational MCTS lies in the expansion phase, as illustrated in Figure 3. In the expansion phase of MCTS, the next move is typically selected from the set of children of the current node. For instance, in MCTS, the candidate moves from node $t1$ are $t3$ and $t4$. In contrast, in Relational MCTS, the candidate moves from node $t1$ include $t3$, $t4$, a $Stop$ action, and the sibling nodes of $t1$. At the same depth level, if two sibling nodes exist, such as in Step 4 of Relational MCTS, the next move is selected from the children of these sibling nodes. For example, when node $t4$ is selected, its candidate moves are $t5$, $t6$, and the $Stop$ action. The inclusion of the $Stop$ action allows Relational MCTS to automatically halt and generate a relevant relational path (1 or 2 hops) for the input query.
RiTeK-PharmKG RiTeK-ADint
Exact Match Rouge-1 Exact Match Rouge-1
Approach P R F1 P R F1 P R F1 P R F1
Zero-Shot GPT-4 11.39 10.90 11.03 15.56 15.50 15.30 7.26 12.10 8.03 13.71 27.64 16.35
+COT Wei et al. (2022) 13.11 16.42 13.70 17.53 22.57 18.40 10.52 19.78 11.95 17.79 37.25 20.97
+Random Walk 12.27 11.86 11.96 14.69 14.15 14.30 15.12 22.68 16.52 20.87 32.92 23.25
+MCTS  Chaslot (2010) 17.17 16.54 16.68 19.09 18.44 18.60 16.97 24.41 18.35 22.82 34.69 25.20
+Relational MCTS (ours) 17.39 20.06 17.57 20.50 23.90 20.82 17.45 24.62 18.76 23.08 34.59 25.27
Few-Shot GPT-4 13.75 15.54 14.04 16.84 19.84 17.49 17.57 17.91 17.48 25.50 28.08 26.04
+COT Wei et al. (2022) 17.29 16.91 16.99 21.55 20.97 21.13 18.57 18.12 18.26 26.68 26.62 26.53
+Random Walk 11.02 13.28 11.32 14.46 17.88 14.92 22.99 22.79 22.75 29.10 29.07 28.95
+MCTS Chaslot (2010) 17.79 17.11 17.30 20.97 20.29 20.48 19.51 27.32 20.91 24.71 36.25 26.96
+Relational MCTS (ours) 19.82 19.03 19.21 22.74 21.81 22.06 23.69 23.72 23.48 30.09 30.36 29.98
@@ 5.2 MCTS vs Relational MCTS
Table 5: Results of various approaches for question answering with complex reasoning on RiTeK-PharmKG and RiTeK-ADint
@@ 6.1 Retrieval Models and Evaluation Metrics
We evaluated five different models on our benchmark datasets, including:
@@
  • Chain-of-Thought Wei et al. (2022): We designed the instruction to guide GPT-4 in generating the answer step by step, with the output formatted as step-by-step reasoning: explanation, answer: medical terms.
  • Random Walk: Starting from the topic entity, a random walk algorithm is applied to explore paths in the textual triple graph in the maximum depth $d$.
  • MCTS Chaslot et al. (2008): Starting from the topic entity, a monte carlo tree search algorithm is applied to explore paths in the textual triple graph in the maximum depth $d$.
  • Relational MCTS.
@@ 6.2 Results and Discussion
Table 5 shows the experiment results of various approaches based on Excat Match and Rougle-1. We have the following observations. (1) We observed that the baseline models struggle to generate the correct answers on RiTeK-PharmKG and RiTeK-ADint. For GPT-4 and GPT+COT, they are challenges in utilizing reasoning information from the graph. Although GPT+COT can utilize step-by-step reasoning, it still relies on the inherent knowledge of the LLM, which limits its ability to apply clear logical reasoning based on knowledge graphs. For the Random Walk, while it can provide reasoning paths, its random nature limits its ability to accurately identify the correct path information. However, it could get the better performance than GPT-4 in RiTek-Adint. Both MCTS and Relational MCTS still face challenges in consistently answering questions and providing the correct reasoning path in 100 percent. However, they can offer valuable path information that enhances LLM performance, especially for Relational LLMs. (2) Our Relational MCTS outperforms all the strong baselines across all evaluation metrics. This success is attributed to the designed stop sign and expansion progress, which help the model identify the appropriate path information effectively. (3) Relational MCTS demonstrates better precision but lower recall in the RiTeK-ADint dataset. This discrepancy arises from the fact that MCTS generates an average of 2.70 answers for each sentence, while RMCTS generates an average of only 1.08 answers. The ground truth average length of answers is 1.02. In the medical domain, we aim to generate more accurate answers.
6.3 The Impact of Hyper-Parameters
In our proposed model, there are two parameters controlling the Relational MCTS: (1) $\epsilon\in\{3000,5000,7000,9000,11000\}$ which is the iteration refers to one complete cycle of the algorithm, where a path is selected, expanded, simulated, and updated to improve decision. (2) $d\in{2,3}$ is the maximum depth of path retrieval. We analyze the impact of these two hyper-parameters in Figure 4 and Figure 5. As we can see, when $\epsilon=5000$, relational MCTS selects paths that are more relevant to the input query and consistently provides better performance than the other settings. Otherwise, we noticed that the model achieves similar performance with $\epsilon=5000$ and $\epsilon=9000$ on the RiTeK-ADint dataset. Since more iterations increase computation time, we ultimately select 5000 iterations. Additionally, $d=3$ outperforms the setting of $d=2$, indicating that when $d=2$, relational paths with more than two hops are not retrieved, which negatively impacts the performance of Relational MCTS. Note: in this section, all experiments are in the zero-shot setting.
Figure 4: The extract match F1 performance of Relational MCTS under different iteration settings.
Figure 5: The extract match F1 performance of Relatinoal MCTS under different hop depth setting
6.4 Case Study
@@
Finally, we highlight the importance of relational path information retrieval ability in achieving good results on our benchmark dataset. In Figure 6, we conduct a case study to compare the performance of Chain-of-Thought, MCTS and Relational MCTS, they are also based on GPT-4. As shown in Figure 6 (green box), without relational path information, Chain-of-Thought failed to generate the right answer Myocardial Reperfusion Injury. Furthermore, MCTS struggles to retrieve the correct path information, resulting in the retrieval of wrong paths that lead the LLM to generate incorrect entities such as Small RNA and MK2206. In contrast, the relational path retrieved by RMCTS could help LLM find the right answer and generate the relevant reasoning information. Specifically, its reasoning explanation information The path then indicates that the interactions of Small RNA lead to dysfunction in organism function which contributes to the pathologic function of damaged myocardium during reperfusion injury is relevant to the final answers. While in COT, the generated explanation is Ultimately, these changes can lead to conditions like heart failure, which is a pathologic function associated with damaged myocardium. Therefore, the final answer is heart failure, which is not relevant to the final answers. In Figure 6 (red box), Relation MCTS and MCTS both failed to get the reasonable path information for the more complex query. They all failed to get the right relational path set by step. This also highlights the challenges present in our benchmark dataset.

### v3

@@ 4.2.3 Data Distribution Analysis
We chose Shannon Entropy and Type-Token Ratio (TTR) as metrics to evaluate query diversity generated in our two datasets. Shannon Entropy takes into account the frequency of each word, measuring the evenness of word distribution in the text, while Type-Token Ratio reflects the variety of words, with a higher value indicating greater diversity in the generated queries. We found the TTR values for both RiTeK-PharmKG and RiTeK-ADint surpass those of STaRK-Prime, demonstrating that the queries generated in our datasets exhibit high complexity and diversity (The results are shown in Appendix D and Table 6). For Shannon Entropy, our results are comparable to STaRK-Prime. Since our RiTeK-ADint dataset involves a wide range of non-pharmacological interventions (NPIs), lifestyle modifications, and environmental factors, it introduces a richer variety of specialized terminology and concepts into the synthesized queries. This expanded vocabulary diversity results in significantly higher Shannon Entropy compared to other medical domain datasets. However, since our two datasets are derived from the medical domain, the frequent repetition of specialized medical terminology, as well as the more concentrated vocabulary compared to general-domain texts, results in slightly lower Shannon Entropy for our datasets than for the other two general-domain datasets. For further analysis about the distribution of query lengths and answer length, please refer Appendix A.6.
RiTeK-PharmKG RiTeK-ADint STaRK-Prime
Exact Match Rouge-1 Exact Match Rouge-1 Exact Match Rouge-1
Approach P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Zero-Shot GPT-4 11.39 10.90 11.03 15.56 15.50 15.30 7.26 12.10 8.03 13.71 27.64 16.35 5.23 6.81 4.65 11.31 16.35 11.31
+Random Walk Lovász (1993) 12.27 11.86 11.96 14.69 14.15 14.30 15.12 22.68 16.52 20.87 32.92 23.25 7.50 8.20 6.48 13.90 17.31 13.32
+MCTS  Chaslot (2010) 17.17 16.54 16.68 19.09 18.44 18.60 16.97 24.41 18.35 22.82 34.69 25.20 7.64 8.36 6.52 14.04 17.45 13.38
+COT Wei et al. (2022) 13.11 16.42 13.70 17.53 22.57 18.40 10.52 19.78 11.95 17.79 37.25 20.97 6.47 8.23 5.81 12.61 17.99 12.47
+TOT Yao et al. (2024a) 7.31 7.32 7.22 13.21 14.67 13.42 3.97 9.65 5.28 12.90 25.44 15.96 2.99 3.08 2.55 9.50 9.81 8.65
+GOT Besta et al. (2024) 3.56 4.20 3.75 10.86 11.84 11.06 2.61 3.32 2.81 15.09 17.63 15.84 1.99 2.20 1.78 9.89 9.34 8.72
+ TOG Sun et al. (2023) 29.85 38.19 31.14 31.38 40.37 32.92 23.08 40.63 25.81 27.81 48.93 31.54 12.14 15.76 11.27 18.67 24.75 18.42
+G-retriever He et al. (2024) 11.21 13.39 11.60 15.01 18.54 15.62 10.97 19.05 12.52 17.27 32.99 20.41 6.23 6.61 5.17 12.01 14.92 11.40
+KAR Xia et al. (2024) 30.95 23.99 25.18 33.65 26.11 27.50 39.59 24.00 27.29 46.54 28.87 32.80 12.02 14.49 11.12 18.04 22.20 17.61
Few-Shot GPT-4 13.75 15.54 14.04 16.84 19.84 17.49 17.57 17.91 17.48 25.50 28.08 26.04 7.79 6.41 5.91 14.03 13.53 12.14
+Random Walk Lovász (1993) 11.02 13.28 11.32 14.46 17.88 14.92 22.99 22.79 22.75 29.10 29.07 28.95 9.93 6.93 7.34 16.54 13.02 13.45
+MCTS Chaslot (2010) 17.79 17.11 17.30 20.97 20.29 20.48 19.51 27.32 20.91 24.71 36.25 26.96 9.57 6.89 7.14 15.92 12.55 12.88
+COT Wei et al. (2022) 17.29 16.91 16.99 21.55 20.97 21.13 18.57 18.12 18.26 26.68 26.62 26.53 8.13 5.91 5.99 14.03 13.53 12.14
+TOT Yao et al. (2024a) 14.74 14.74 14.63 19.22 19.14 18.97 13.28 13.17 13.21 24.65 24.72 24.60 12.84 10.11 10.36 6.93 4.85 5.06
+GOT Besta et al. (2024) 12.10 12.22 12.06 17.38 17.31 17.19 15.84 15.32 15.42 26.20 25.89 25.91 5.37 3.73 3.78 12.69 9.98 10.17
+TOG Sun et al. (2023) 29.14 42.33 32.36 30.40 44.00 33.88 26.50 47.13 33.83 29.46 49.69 36.43 14.41 20.39 16.40 19.75 26.61 20.14
+G-retriever He et al. (2024) 12.51 12.14 12.22 15.94 15.44 15.57 17.47 17.50 17.32 24.87 24.92 24.71 7.72 5.75 5.86 14.63 11.92 12.10
+KAR Xia et al. (2024) 27.35 27.43 26.99 29.74 29.76 29.34 34.68 33.42 33.48 40.15 38.55 38.88 13.01 15.50 12.21 19.00 23.10 18.00
Supervised G-retriever He et al. (2024) 38.71 37.11 37.62 39.78 39.18 39.31 47.93 47.16 47.41 54.68 54.00 54.24 16.14 16.47 14.11 17.21 27.86 19.21
GCR Luo et al. (2024) 44.38 57.28 47.71 46.04 58.83 49.44 43.52 60.78 48.07 49.47 65.57 54.24 19.03 26.89 18.94 28.01 37.18 28.75
GNN-RAG Mavromatis and Karypis (2024) 50.78 49.28 49.72 51.66 50.29 50.73 51.04 50.59 50.55 56.49 56.09 56.09 16.00 15.04 14.50 24.78 23.51 22.99
@@
Table 4: Results of various approaches for question answering with complex reasoning on RiTeK-PharmKG, RiTeK-ADint and STaRK-Prime. P refers to the Precision, R refers to the recall. In the experiments, the GPT-4 version is GPT4o-mini.
@@ 5.1 Retrieval Models and Evaluation Metrics
We evaluated 11 representative retrieval models on our benchmark datasets under both zero-shot and few-shot learning settings. In addition to our benchmark dataset, we also evaluated the models on STaRK-Prime Wu et al. (2024b), a textual question answering dataset with minimal ontological structure in its queries, including:
@@
  • Random Walk Lovász (1993): Starting from the topic entity, a random walk algorithm is applied to explore paths in the textual triple graph in the maximum depth $d$.
  • MCTS Chaslot (2010): Starting from the topic entity, a Monte Carlo tree search algorithm is applied to explore paths in the textual triple graph in the maximum depth $d$. In this work, we set the $d=3$.
  • Chain-of-Thought (COT) Wei et al. (2022): We designed the instruction to guide GPT-4 in generating the answer step by step, with the output formatted as step-by-step reasoning: explanation, answer: medical terms.
  • Tree-of-Thought (TOT) Yao et al. (2024a):We structured the reasoning process as a tree search, where multiple intermediate reasoning paths are explored in parallel. GPT-4 evaluates and expands promising paths based on a voting or scoring mechanism.
  • Graph-of-Thought (GOT) Besta et al. (2024): We represented the reasoning process as a graph structure, where nodes capture different reasoning states and edges denote transitions. GPT-4 traverses the graph to aggregate information and synthesize the final answer.
  • Think-on-Graph (TOG) Sun et al. (2023): is a reasoning framework that enables large language models to interactively perform beam search over knowledge graphs, discovering and evaluating promising reasoning paths without additional training.
  • G-retriever He et al. (2024): A RAG-based approach that retrieves query-relevant subgraphs using the Prize-Collecting Steiner Tree (PCST) algorithm to enhance LLM understanding and reasoning over textual graphs.
  • KAR Xia et al. (2024): A knowledge-aware query expansion method that augments LLMs with structured document relations from a knowledge graph, using relation-aware filtering to improve retrieval for semi-structured queries.
We evaluated the 3 representative retrieval models on our benchmark datasets and STaRK-Prime under supervised learning settings, including:
  • G-retriever He et al. (2024):A RAG-based approach that retrieves query-relevant subgraphs using the Prize-Collecting Steiner Tree (PCST) algorithm to enhance LLM understanding and reasoning over textual graphs.
  • GCR Luo et al. (2024): A knowledge-aware query expansion method that augments LLMs with document-based relational signals to improve retrieval for semi-structured queries.
  • GNN-RAG Mavromatis and Karypis (2024). A method that uses a GNN to retrieve relevant answers and extract the shortest paths connecting the topic entity and answers, which are then verbalized and fed into the LLM to enhance retrieval-augmented generation (RAG) performance.
@@ 5.2 Results and Discussion
Table 4 shows the experiment results of various approaches based on Exact Match and Rouge-1. We have the following observations. Zero-shot and few-shot setting: (1) We observed that the baseline models struggle to generate the correct answers on RiTeK-PharmKG and RiTeK-ADint. For GPT-4 and GPT+COT, they are challenges in utilizing reasoning information from the graph. Although GPT+COT can utilize step-by-step reasoning, it still relies on the inherent knowledge of the LLM, which limits its ability to apply clear logical reasoning based on knowledge graphs. For the Random Walk, while it can provide reasoning paths, its random nature limits its ability to accurately identify the correct path information. However, it could get the better performance than GPT-4 in RiTek-Adint and STaRK-Prime in the zero/few-shot setting. (2) Tree-of-Thought (ToT) and Graph-of-Thought (GoT) attempt to guide LLM reasoning through structured prompting, encouraging step-by-step or graph-based logical thinking. However, on complex textual KBQA datasets like RiTeK-PharmKG and RiTeK-ADint, both methods consistently underperform, with F1 scores far below those of retrieval-augmented approaches like KAR (e.g., ToT: 13.42 vs. KAR: 27.50 in zero-shot). This suggests that the internal knowledge and reasoning capabilities of LLMs alone are insufficient for tasks that require fine-grained relational understanding and the integration of attribute information from the query. Despite their logical scaffolding, ToT and GoT struggle to recover factual precision without access to external structured knowledge. (3) KAR achieves strong performance on medical datasets like RiTeK-PharmKG and RiTeK-ADint, outperforming baselines in both zero-shot and few-shot settings. Its main strength lies in combining textual semantics with structured KG relations to generate accurate and context-aware query expansions. However, KAR relies on retrieving the top-n relevant documents; however, determining an appropriate value for n and the optimal order in which to select documents is non-trivial. (4) G-Retriever shows moderate performance across medical datasets, but generally underperforms compared to methods like KAR or TOG in both zero-shot and few-shot settings. For example, on STaRK-Prime, its ROUGE-1 F1 score (5.17 vs. 11.12 zero-shot) lags significantly behind KAR, This indicates a weaker ability to handle complex relational constraints, particularly when the answer’s attributes are embedded in the query. Its main strength lies in interpretable subgraph selection using PCST, which enhances explainability and helps mitigate hallucinations. (5)TOG performs moderately in zero-shot settings but shows strong gains in few-shot scenarios, achieving top-tier ROUGE-1 F1 scores like 37.11 on RiTeK-ADint and 36.43 on STaRK-Prime. This highlights its ability to leverage demonstrations to guide accurate reasoning over knowledge graphs, especially in complex biomedical tasks.
In the setting of supervised fine tuning, GCR achieves the best overall performance across all three medical benchmarks in the supervised setting, with scores like 57.28 ROUGE-1 F1 on ADint and 49.72 on STaRK-Prime, demonstrating its strength in generating faithful, KG-grounded answers. However, GCR relies on pre-constructed KG-Trie indices. We found that GNN-RAG achieves better performance on the RiTeK-PharmKG and RiTeK-ADint datasets, demonstrating its ability to retrieve relevant path information from the graph. However, since it primarily relies on shortest paths, it may overlook critical reasoning information embedded in more complex or indirect graph structures.
llama 3.1 8b llama2-chat-7b Biomixtral 7b
Exact Match Rouge-1 Exact Match Rouge-1 Exact Match Rouge-1
Approach P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
RiTeK-PharmKG G-retriver He et al. (2024) 36.97 46.07 38.31 33.04 47.03 38.41 38.71 37.11 37.62 39.78 39.18 39.31 43.01 41.59 42.01 43.95 42.69 43.10
GNN-RAG 33.21 43.01 37.31 21.00 44.89 26.00 50.78 49.28 49.72 51.66 50.29 50.73 39.93 39.12 39.26 41.69 40.89 41.08
w/o retriever 32.45 43.40 34.23 47.60 46.59 46.84 38.91 37.63 38.02 40.57 39.31 39.72 41.49 39.43 40.05 41.25 41.06 40.99
RiTeK-ADint G-retriver He et al. (2024) 50.83 50.07 50.31 57.34 56.61 56.87 47.93 47.16 47.41 54.68 54.00 54.24 48.34 47.48 47.75 55.58 54.73 55.02
GNN-RAG 40.88 40.90 40.43 44.43 45.01 45.49 51.04 50.59 50.55 56.49 56.09 56.09 50.83 50.07 50.31 57.34 56.61 56.87
w/o retriever 49.59 48.48 48.82 55.23 54.29 54.61 46.58 45.82 46.06 51.66 49.91 46.47 49.79 48.93 49.20 56.43 53.80 54.15
STaRK-Prime G-retriver He et al. (2024) 16.14 16.47 14.11 17.21 27.86 19.21 10.15 8.45 8.17 21.75 18.08 18.40 12.22 11.54 10.58 23.15 21.37 20.72
GNN-RAG 7.81 16.67 9.35 18.13 27.50 19.65 16.00 15.04 14.50 24.78 23.51 22.99 11.20 10.31 10.65 17.98 18.09 18.32
w/o retriever 12.96 14.99 11.91 16.80 25.75 18.41 11.77 10.38 9.68 20.65 21.59 17.66 12.96 11.12 10.73 24.83 21.59 21.68
Table 5: Performance of different retrieval models across backbone LLMs. “w/o retriever” denotes an LLM without retrieval augmentation.
5.3 Analysis
5.3.1 Effect of Different LLMs on Retriever Effectiveness
In this part, we analyze the influence of different LLMs on the retrievers. Table 5 presents the performance of three retrieval settings, G-retriever, GNN-RAG, and without retriever, in three LLMs of the backbone: Llama 3.1 8b, Llama2-chat-7b, and Biomixtral 7b, on three datasets. Overall, G-retriever consistently outperforms other approaches across most metrics, particularly in Rouge-1 F1 scores. For instance, on RiTeK-ADint, G-retriever with Llama 3.1 8b achieves the highest F1 score of 56.87, while the GNN-RAG and no retriever baselines lag behind. Similarly, G-retriever reaches 55.02 F1 on Biomixtral for the same dataset, showcasing its robustness across model sizes. In contrast, GNN-RAG shows variable performance, sometimes underperforming even compared to the no-retriever baseline, such as on STaRK-Prime using Biomixtral. The "w/o retriever" baseline, representing an LLM without retrieval augmentation, performs surprisingly well in some settings, indicating that strong LLMs alone can capture a significant amount of relevant knowledge. For example, on RiTeK-ADint with Biomixtral, it achieves a Rouge-1 F1 score of 54.15, close to the GNN-RAG. However, in most cases, retrieval-augmented methods still yield superior performance. Notably, Biomixtral 7b tends to outperform the other two LLMs when combined with retrieval, especially in recall and F1. These results suggest that both the choice of retrieval strategy and the backbone LLM significantly impact end-task performance.
5.3.2 Case Study of Path and Answer Quality
@@
We conduct a qualitative analysis to compare the reasoning paths and predicted answers from different retrieval models on two biomedical question-answering examples. As shown in Figure 2, all models successfully predicted the correct answer Myocardial Reperfusion Injury in the first case, although their reasoning paths varied in granularity and relevance. G-retriever and GNN-RAG produced informative multi-hop paths that partially overlapped with the ground truth.
In contrast, for the second question involving CHI3L1 and schizophrenia, only the ground truth path led to the correct answer. All baseline models failed: TOG and MCTS generated incorrect reasoning chains centered around Alzheimer’s Disease and Juvenile Huntington Disease, while GPT-4 hallucinated Autism Spectrum Disorder. These errors reveal the challenge of modeling rare or indirect biomedical associations, especially when entity relations involve subtle phenotypic markers. This case highlights the importance of precise multi-hop reasoning and clinically aligned retrieval in semi-structured biomedical graphs.